Circular No. 5935

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

PERIODIC COMET HALLEY (1986 III) The following positions have been derived from observations made by

O. Hainaut (in a joint program with R. M. West, B. G. Marsden and K. J. Meech) with the 3.5-m New Technology Telescope at the European Southern Observatory. Nine 25-min exposures in succession with the SuSI CCD camera and V filter, tracked at the motion of the comet, on Jan. 11 show an image at V = 26.5 ± 0.2 in all frames (combining to a 7σ signal); seeing was 0′′.6–0′′.9 and the sky surface brightness V = 21.9 mag arcsec^{−2}. The positions, from three trios of images, show that the comet’s motion is well within 5 percent of what was expected, whereas a main-belt minor planet in the relevant part of the sky would be expected to move thrice as fast (and a Trojan twice as fast). At a heliocentric distance of 18.8 AU, the comet seems to be the faintest ever recorded. These observations confirm the probable presence of an unmeasurable 2σ image near the limit of seven 10-min and four 15-min exposures on 1992 Apr. 6 (cf. IAUC 5535), when the seeing was 1′′.2 and the sky background V = 20.8 mag arcsec^{−2}.

1994 UT  α_{2000}  δ_{2000}
Jan. 11.19243  9^{h}15^{m}03.^{s}68  −3^{o}21′59′′.7
11.25655  9 15 02.92  −3 21 57.8
11.32237  9 15 02.09  −3 21 56.2

PERIODIC COMET KUSHIDA (1994a)

Orbital elements from MPC 23105:

T = 1993 Dec. 12.8428 TT  ω = 214.^{o}4507
e = 0.639349  Ω = 245.9597 }2000.0
q = 1.367357 AU  i = 4.1863

a = 3.791356 AU n^{o} = 0.1335095 P = 7.382 years

1994 TT α_{2000}  δ_{2000}  Δ  r  ϵ  β  m_{1}
Feb. 27 9^{h}41.^{m}73 + 2^{o}38.^{′}9 0.647 1.623 164.^{o}5 9.^{o}4 12.2
Mar. 9 9 41.93 + 3 30.2 0.726 1.684 156.3 13.7 12.6
19 9 44.75 + 4 13.0 0.824 1.748 147.6 17.7 13.0
29 9 50.13 + 4 41.5 0.937 1.815 139.3 21.0 13.4
Apr. 8 9 57.69 + 4 53.9 1.066 1.884 131.4 23.5 13.9
18 10 07.03 + 4 50.4 1.208 1.955 124.0 25.2 14.3
28 10 17.75 + 4 32.1 1.362 2.027 116.9 26.3 14.7
May 8 10 29.48 + 4 01.1 1.525 2.100 110.2 26.8 15.1

1994 February 18 Brian G. Marsden